OpenStudy (the_puppet1):

Solve the equation 2q - 4 = 26 for q. A. 15 B. 20 C. 25 D. 30

OpenStudy (ifrah123):

15

OpenStudy (carolina4567):

Answer

OpenStudy (dannyrod2000):

Simplifying 2q + -4 = 26 Reorder the terms: -4 + 2q = 26 Solving -4 + 2q = 26 Solving for variable 'q'. Move all terms containing q to the left, all other terms to the right. Add '4' to each side of the equation. -4 + 4 + 2q = 26 + 4 Combine like terms: -4 + 4 = 0 0 + 2q = 26 + 4 2q = 26 + 4 Combine like terms: 26 + 4 = 30 2q = 30 Divide each side by '2'. q = 15 Simplifying q = 15

OpenStudy (the_puppet1):

I think it's A

OpenStudy (carolina4567):

Good Job